An animal takes a breath with a volume of 1.80 L at standard temperature and pressure. If air contains a 20.9% oxygen, how many molecules of oxygen are in one breath?
1
Expert's answer
2019-10-18T08:07:02-0400

Air inhaled"=1.80 \ L"

Oxygen inhaled"=20.9" % of air"=\frac{20.9}{100}\times 1.80=0.3762 \ L"

Number of moles of "O_2" in this amount"=\frac{0.3762}{22.4}=16.795\times 10^{-3} \ mol"

Number of molecules of "O_2=16.795\times10^{-3}\times 6.023 \times 10^{23}=101.156\times10^{20}"